# Redefining Toxicity: An Objective and Context-Aware Approach for Stress-Level-Based Detection

### v2

@@ 4.1 Towards Measurement: Stress as an Objective Metric
Non-invasive biosignals, including heart rate variability (HRV), galvanic skin response (GSR), pupil dilation, respiration, and facial or vocal affect, provide scalable approximations of stress. These signals correlate reliably with social and emotional threat responses Sharma and Gedeon (2012); Giannakakis et al. (2019), but still depend on hardware access and stable observation conditions.
Behavioural proxies, such as language use, reaction patterns, and discourse dynamics, offer the most feasible solution for large-scale deployment. In online settings, sentiment-laden replies, conversational shifts, and reaction memes often serve as collective indicators of stress. Prior research shows that these patterns reliably signal perceived norm violation and emotional disruption Jafari et al. (2023); Aleksandric et al. (2024).
5 PONOS: A Metric for Perceived Toxicity
Accepting toxicity being reflected in the evoked stress, we obtain observable proxy for stress levels. When a post elicits a wave of negative comments, disagreement, or distress, we see this as an indicator of perceived toxicity.
@@ 5 PONOS: A Metric for Perceived Toxicity
To formalise this, we introduce the Proportion Of Negative Observed Sentiments, abbreviated as PONOS (from Greek πόνος - "pain") — as an instantiation of our stress-based model. PONOS quantifies toxicity not through content analysis or classifier predictions, but through the emotional reception of a message within a specific community.
@@ 5.1 Defining PONOS
Let a post $x$ receive a set of replies $R(x)=\{r_{1},r_{2},\dots,r_{n}\}$. Each reply $r_{i}\in R(x)$ is assigned a sentiment score $s(r_{i})\in\{-1,0,+1\}$, where $-1$ denotes negative sentiment, 00 is neutral, and $+1$ is positive.
@@
$\text{PONOS}(x)=\frac{1}{|R(x)|}\sum_{r_{i}\in R(x)}\mathbb{I}[s(r_{i})=-1]$ (1)
where $\mathbb{I}[\cdot]$ is the indicator function.
This metric quantifies the perceived negativity of a post based on its community response. Higher values of $\text{PONOS}(x)$ indicate a stronger collective expression of disapproval or tension.
Compared to conventional approaches such as lexical sentiment counts, softmax-based toxicity classifiers, or black-box ML scores, PONOS metric offers the following advantages:
  • Interpretability: The value of PONOS directly reflects the proportion of negative replies, making it transparent and accessible to users, moderators, and downstream systems.
  • Context-relativeness: Because it reflects local community reactions, PONOS adapts to group norms, dialects, humour, and platform-specific conventions without requiring explicit encoding.
  • Statistical grounding: PONOS can be understood as the empirical mean of a Bernoulli distribution over negative/non-negative labels. By the law of large numbers, it converges to the true community-level disapproval rate as the number of replies increases — unlike softmax-based model scores, which often lack a consistent probabilistic interpretation (Guo et al., 2017).
  • Ratio-scale comparability: Because the score is ratio-scaled, posts with higher PONOS values are consistently more negatively received. For example, a score of 0.6 indicates twice the disapproval rate of 0.3. This enables meaningful ranking, thresholding, and aggregation — unlike many opaque ML toxicity scores, which cannot be meaningfully compared across models.
PONOS is a flexible framework that admits multiple extensions to better reflect community dynamics, reply timing, and user context. In Appendix B we describe its variants, such as PONOS-Net, PONOS-Weighted, and PONOS-Early.
PONOS is not a universal measure of toxicity. It is a context-dependent metric — a community-grounded, reply-driven proxy for perceived social tension. Rather than determining what a message intrinsically "is", PONOS captures how it is received - reflecting the social impact of communication in its local context.
@@ 6 Empirical Evaluation
To demonstrate the viability of our framework for context-sensitive toxicity detection, we conducted a targeted empirical study. The process involves two main steps: first, identifying the most suitable sentiment function $s(\cdot)$ for the given context, and second, using it to select the best reaction prediction model for obtaining the PONOS value.
@@ 6.1 Dataset Creation
We curated a dataset from the Reddit community r/BlackPeopleTwitter. This community was chosen due to its large size (5.8 million members) and the fact that its guidelines explicitly encourage the use of African-American Vernacular English (AAVE) and slang.
Examples of AAVE often get misclassified as toxic speech by current toxicity detection models, which generally ignore contextual and cultural nuances. For example, a supportive phrase in AAVE like "That’s my nigga!" (roughly equivalent to “That’s my man!” in Standard American English) might be wrongly labelled as hateful or toxic. Such misclassifications effectively penalise speakers for communicating in their own dialect.
@@
6.1.1 Annotation Process
To identify the optimal sentiment function $s(\cdot)$, we created a human-labeled test set from a subset of the data. We selected 56 user messages, each with at least four replies. If a message had too many replies, we retained only the top 7 by score, resulting in a total of 273 replies.
We conducted a structured labelling task using the Prolific platform. Annotators evaluated the sentiment of each reply, assigning two types of labels: (1) sentiment polarity — approval, neutral, or condemnation, and (2) sentiment target — whether the sentiment was directed at the original comment or the topic being discussed. This produced compound labels such as “approval (comment)” or “condemnation (subject).”
We recruited 11 annotators, all self-identifying as Black/African American U.S. nationals and regular internet users. The group included 7 men and 4 women, aged 21 to 60. To ensure high-quality data, we embedded attention checks and analysed labelling consistency. Four annotators showed strong engagement and reliable annotations; their labels were retained for further analysis.
We assessed inter-annotator agreement, obtaining a Fleiss’ Kappa of 0.48 and pairwise agreement of 67.8%, indicating moderately strong consensus given the subjectivity of the task. Final labels were determined using majority vote. In rare cases of a 2–2 tie, an additional annotation was provided by a researcher using the same labelling guidelines.
6.1.2 Ethical Considerations
To ensure our study adhered to ethical research standards, we took several measures. First, we complied with Reddit’s terms of service and ensured that no sensitive personal data was collected. We explained the nature of the task and data collection process to the participants beforehand, ensuring they were fully aware that the collected data would be used for research purposes.
Participants were compensated at a rate of \$12 per hour, in accordance with Prolific’s recommended payment guidelines. To estimate the time required for survey completion, we first completed the survey ourselves. The full text of the instructions given to participants can be found in Appendix D.
6.2 Experiments
We conducted all experiments using NVIDIA H100 GPUs, totaling 883 compute hours. Model training and evaluation were performed using the February 2025 release of the Unsloth library (unslothai, 2025). Metrics were computed using scikit-learn version 1.5.0 Pedregosa et al. (2011).
We were committed to providing as much context as possible to the models, and we found this to be beneficial for the models’ performance. The input data for sentiment classification of a reply to a comment included: the comment itself, the text of all replies for classification, the title of the post. The prompt used with LLMs is provided in Appendix E.
6.2.1 Sentiment Analysis Model
To determine the most suitable sentiment analysis function $s(\cdot)$ for use in Equation 1, we evaluated both specialized sentiment classifiers and general-purpose LLMs using custom prompts. Evaluation was conducted on the human-labelled subset of our dataset. Results are presented in Table 1.
@@ 6.2.1 Sentiment Analysis Model
RoBERTa (2023) 0.63 0.22 0.32
ModernBERT (2024) 0.71 0.21 0.32
Twitter-RoBERTa (2020) 0.58 0.64 0.61
GPT-4o (2024a) 0.67 0.82 0.74
@@
Table 1: Precision, recall, and F1 scores for negative sentiment class across various models.
@@
Qwen 3 Yang et al. (2025) showed the best overall performance and was selected for all subsequent experiments. During error analysis, we found that most misclassifications occurred in examples with low inter-annotator agreement. These instances often involved ambiguous language that led both assessors and models to varied interpretations.
6.2.2 Toxicity Prediction
With a sentiment analysis model selected, we computed PONOS scores for each post in the dataset. To evaluate binary toxicity classification, we thresholded the continuous PONOS values: posts with $\text{PONOS}\geq 0.5$ were labeled as toxic, and all others as non-toxic. The threshold of 0.5 reflects the intuition that at least half of readers are likely to respond negatively. It also aligns with thresholds used in many existing moderation frameworks, facilitating fair comparison with baseline models.
We made the following evaluations:
Direct PONOS Prediction:
A regression model was trained to predict the gold-standard PONOS score from post text alone. This method is simple and computationally efficient, making it suitable for real-time toxicity detection using lightweight models.
PONOS Prediction via Simulated Reactions:
A language model was trained to generate plausible replies to a given post. These synthetic replies were passed through the sentiment analysis model to estimate the PONOS score. This method aligns with our theoretical framing of toxicity as an emergent social signal: it approximates how a community might react.
Comparison with Baselines:
As baselines, we include predictions from the OpenAI Moderation API (omni-moderation-latest) (2024b), Google Perspective API (2024), and Detoxify (2020). These systems represent current industry standards and enable comparison in terms of contextual sensitivity and error types.
Ablation Study:
To isolate the contribution of domain-specific adaptation and reaction-modeling strategies, we conducted an ablation study using LLaMA 3 70B Llama Team (2024) across four variants:
  • TP (Toxicity Prediction): Predicts binary toxicity label directly from the input utterance.
  • DTTP (Domain-Tuned Toxicity Prediction): Same as TP, but with domain-specific pretraining on the unlabeled portion of our dataset.
  • RP (Reply Prediction): Generates replies to the input. Toxicity is then inferred by applying a thresholded PONOS score computed from the predicted replies using a sentiment analysis model.
  • DTRP (Domain-Tuned Reply Prediction): Same as RP, but with domain-specific pretraining on the unlabeled dataset.
6.2.3 Large Language Model Pretraining
In DTTP and DTRP variants we pretrained a LLaMA 3 70B model using domain-specific Reddit data from r/BlackPeopleTwitter. The objective was to perform domain adaptation, enabling the model to generate plausible, community-authentic replies (DTRP) or contextually-aware toxicity prediction (DTTP).
Model Architecture and Configuration
@@ 6.2.3 Large Language Model Pretraining
We used the unsloth framework to load and fine-tune the base model in 4-bit precision. The model was adapted using LoRA with rank 64 for efficient parameter updates. Gradient checkpointing and rslora were employed to further reduce memory overhead. All training and inference were conducted using bfloat16 precision and FlashAttention 2 to maximize memory efficiency.
Prompt Format and Dataset Construction
@@
Each training example consisted of a Reddit comment followed by a prompt requesting five likely replies in the tone and style of r/BlackPeopleTwitter. Replies were separated by a special delimiter token <|reply|>. The prompt template is provided in Appendix E.
@@
Table 2: Mean Squared Error (MSE) and Mean Absolute Error (MAE) for PONOS score prediction . FT = finetuned model.
Model Prec. Rec. F1
Direct Prediction
TF-IDF + Ridge 0.63 0.26 0.36
all-mpnet + Ridge 0.65 0.30 0.41
SentenceBERT FT 0.54 0.65 0.59
Pretrained Baselines
OpenAI moderation 0.43 0.28 0.34
Perspective API 0.35 0.31 0.33
Detoxify 0.35 0.31 0.33
Detoxify-unbiased 0.35 0.34 0.35
Reply Prediction and Ablation Study
LLaMA 3 70B TP 0.43 0.40 0.42
LLaMA 3 70B DTTP 0.42 0.42 0.42
LLaMA 3 70B RP 0.40 0.21 0.27
LLaMA 3 70B DTRP 0.67 0.71 0.69
Table 3: Precision, recall, and F1 score for toxicity classification (toxic class only) across direct prediction, pretrained baseline, and reply-based models.
6.3 Results
As shown in Tables 3 and 2, our best model - DTRP (Domain-Tuned Reply Prediction), achieves a substantial improvement in F1 score (0.69), significantly outperforming both direct PONOS predictors (e.g., finetuned SentenceBERT (2019) at 0.59) and external moderation tools such as Detoxify, OpenAI Moderation API, and Perspective API (F1 $\approx$ 0.33–0.35). This confirms the core hypothesis of our work: toxicity is better modelled as contextual social friction than as a fixed textual property.
Ablation Study further illustrates that it’s the combination of domain adaptation and reply-based modelling show the greatest gains.
On top of that, simulated reactions offer several advantages: they incorporate social and cultural context, allow integration of multimodal input (via multimodal LLMs), and provide interpretability by exposing intermediate responses. This approach also generalizes well to novel or underspecified content.
Complementing these quantitative results, our error analysis (Appendix C) reveals the systemic limitations of existing models.
Commercial tools frequently miss ideologically disruptive or passive-aggressive content (false negatives), while overflagging vernacular, humor, and identity expression (false positives). These failures reflect an overreliance on surface features and global thresholds. By contrast, PONOS captures perceived harm through community reactions, allowing our models to align more closely with socially grounded definitions of toxicity.
7 Conclusion: Toward a Contextual Ethics of Harm
Toxicity is not an intrinsic property of language, but an emergent social signal - a reaction to perceived norm violation and induced stress within a particular context. In this framing, harm is not universal but situational; not static but dynamic. What matters is not what is said in isolation, but how it is received and interpreted.
By reframing toxicity detection as the measurement of perceived social friction, we challenge NLP’s prevailing paradigm of lexical essentialism. Instead, we propose a framework grounded in psychology, neuroscience, and computational social science.
Our metric, PONOS, embodies this shift. Rather than classifying messages based on intrinsic properties, PONOS captures their collective social reception. It is interpretable, context-sensitive, and empirically grounded - a pragmatic tool for modelling harm as it is experienced by real communities.
PONOS is not a moral arbiter. It does not tell us what is right, but what is stressful. It is a mirror, not a judge. Moderation systems should treat PONOS as an advisory signal, not a censorship trigger. High PONOS scores do not necessarily indicate ethical failure; they signal social dissonance. Low scores are not always desirable either - total agreement may reflect echo chambers or silencing.
We are not offering a final definition of toxicity, but a new conceptual framework for studying it - one that acknowledges context, perception, and affect. PONOS illustrates that subjectivity can be modelled without being reduced.
Future work can extend this approach by improving the model, incorporating multimodal signals, conducting cross-cultural comparisons, and validating it in real-world scenarios.
Ultimately, we aim to shift NLP from classifying bad words to understanding social dynamics - from linguistic essentialism to social contextualism.

### v3

@@ 4.1 Towards Measurement: Stress as a Measurable Signal
Stress can be estimated through multiple classes of indicators, each trading off fidelity against scalability. Direct physiological measures (e.g. cortisol levels) are biologically faithful but require invasive sampling and are impractical outside controlled settings. Non-invasive biosignals such as heart-rate variability, galvanic skin response, respiration, pupil dilation, and facial or vocal affect provide scalable approximations and correlate with social-threat responses sharma2012objective; soc_stress_biosignals, but still depend on specialised hardware and stable observation conditions.
For NLP systems deployed in online environments, the most practical evidence is therefore behavioural: stress and disruption often surface in interaction through reaction patterns, shifts in tone, escalation, withdrawal, or affective language in replies. Prior work shows that collective response dynamics can serve as indirect indicators of perceived norm violation and emotional disruption in online discussion Amir; Aleksandric_2024. Importantly, behavioural proxies do not measure internal physiology directly; they capture how social tension manifests publicly in interaction - precisely the level at which toxicity becomes consequential for communities and moderation decisions.
5 PONOS: A Metric for Reception-Based Toxicity
Building on the stress-based account of toxicity developed above, we now introduce a concrete metric for capturing reception-based toxicity: how a communicative act is received by its audience in context.
Our definition implies a causal pathway: a post is toxic when it is perceived to contradict local moral or interactional norms, and this perceived violation induces a stress response in the audience.
In online interaction, stress and social disruption often surface as expressions of disapproval, outrage, distress, or condemnation in replies. Accepting that toxicity is reflected in the stress it evokes, we treat patterns of community response as a proxy signal for reception-based toxicity: when a post elicits a disproportionate share of negative reactions, this provides evidence that it violated local norms and induced tension in its audience.
@@ 5 PONOS: A Metric for Reception-Based Toxicity
To formalise this intuition, we introduce PONOS (Proportion Of Negative Observed Sentiment; from Greek πόνος, "pain") as one instantiation of the Contextual Stress Framework. PONOS quantifies toxicity not through lexical content or opaque model scores, but through the emotional reception of a message within a specific community and context.
@@ 5.1 Defining PONOS
Let a post $x$ receive a set of replies $R(x)=\{r_{1},r_{2},\dots,r_{n}\}$. Each reply $r_{i}\in R(x)$ is assigned a sentiment label $s(r_{i})\in\{-1,0,+1\}$, where 1-1 denotes negative sentiment, 0 neutral sentiment, and +1+1 positive sentiment.
@@
$\text{PONOS}(x)=\frac{1}{|R(x)|}\sum_{r_{i}\in R(x)}\mathbb{I}[s(r_{i})=-1],$ (1)
where $\mathbb{I}[\cdot]$ denotes the indicator function.
Intuitively, $\text{PONOS}(x)$ estimates the share of observed replies that express negative affect. Higher values indicate stronger collective expressions of disapproval or distress, consistent with reception-based harm in context.
PONOS is (i) interpretable as a proportion of negative replies, (ii) context-sensitive via local audience norms, and (iii) statistically grounded and comparable, as an estimate of a negative-reception rate whose uncertainty decreases with $|R(x)|$ (guo2017calibration).
PONOS is a flexible framework that admits multiple extensions to better reflect community dynamics, reply timing, and user context. In Appendix C we describe its variants, such as PONOS-Net, PONOS-Weighted, and PONOS-Early.
@@ 6 Empirical Evaluation
This section provides empirical support for CSF: toxicity is not a fixed property of text in isolation, but a contextual social signal that is realised in interaction. Our definition decomposes toxicity into two components: (i) perceived norm violation (contradiction of accepted morality and norms of interaction) and (ii) the stress or disruption this perceived violation induces in an audience. Accordingly, we evaluate PONOS as a reception-based measure of the second component (stress or disruption manifested behaviourally), and we treat widely used text-only toxicity instruments as reference measures that primarily operationalise the first component (perceived norm violation) under assumed or averaged contexts.
We therefore design the evaluation as validity evidence for a reception-based axis, rather than as a leaderboard. Concretely, we test (i) whether the reception signal can be instantiated reliably in-domain, (ii) whether it is estimable without access to future replies, and (iii) whether it provides information that is distinct from intrinsic (text-only) instruments. We operationalise these three goals as the following evaluation questions:
  • Q1 (Operationalisation / instrument selection): Which sentiment model yields an in-domain sentiment function $s(\cdot)$ that best matches human judgements of reply sentiment in this community?
  • Q2 (Estimability at inference time): Can measured PONOS be predicted from the post text alone, and does explicitly modelling likely reactions improve estimation?
  • Q3 (Construct divergence / complementarity): How does reception-based measurement (PONOS) relate to intrinsic-toxicity instruments (text-only toxicity detectors)? Where do they agree, and where do they systematically diverge? Does PONOS provide a complementary axis aligned with CSF, or is it reducible to existing intrinsic scores?
@@ 6.1 Dataset Creation
We curate a dataset from the Reddit community r/BlackPeopleTwitter. We choose this community for two reasons: (i) its scale (5.8M members) and (ii) explicit encouragement of African-American Vernacular English (AAVE) and community-specific slang use. This setting is useful for studying context sensitivity, since dialectal and in-group expressions are frequently misinterpreted by generic toxicity systems trained on broad web data, effectively penalising speakers for communicating in their own dialect.
@@
6.1.1 Human Annotation for Selecting $s(\cdot)$
To instantiate the sentiment function $s(\cdot)$ in Eq. 1, we construct a small, high-quality labelled set of posts and replies, since replies constitute the reception signal from which PONOS is computed.
We sample 56 user messages, each with at least four replies. If a message had too many replies, we retained only the top seven by score, resulting in a total of 273 replies.
Annotators and quality control.
We recruit 11 annotators via Prolific platform, all self-identifying as Black/African American U.S. nationals and regular internet users (7 men, 4 women; ages 21–60). We include attention checks and examine consistency. Four annotators show strong engagement and reliable performance; we retain their annotations for analysis.
Agreement and aggregation.
Inter-annotator agreement is Fleiss’ $\kappa=0.48$ with 67.8% pairwise agreement, consistent with moderate agreement for subjective judgement tasks. Final labels are obtained by majority vote; in rare 2-2 ties, an additional label is provided by a researcher following the same guidelines.
6.2 Experimental Setup
All experiments were run on NVIDIA H100 GPUs, totalling 883 compute hours. Model training and evaluation were performed using the February 2025 release of the unsloth library (unslothai2024). Metrics were computed using scikit-learn v1.5.0 scikit-learn.
Context provided to models.
For reply sentiment classification, we provide as much conversational context as feasible: the original comment, the reply text, the title of the post, and (when applicable) the set of replies under classification. We found that providing this context improved model’s performance. Prompts used with LLMs are provided in Appendix E.
6.3 Step 1 (Q1): In-Domain Instrument Selection for $s(\cdot)$
We evaluate multiple candidate sentiment models on the labelled reply set in order to select $s(\cdot)$ for Eq. 1. This step is best viewed as instrument selection: we choose the model that most reliably reproduces in-domain human judgements of negative reception, rather than optimising for generic sentiment benchmarks.
Table 1 reports performance for detecting the negative sentiment class on our labelled data.
@@ 6.3 Step 1 (Q1): In-Domain Instrument Selection for $s(\cdot)$
RoBERTa (samlowe2023roberta) 0.63 0.22 0.32
ModernBERT (cirimus2024modernbert) 0.71 0.21 0.32
Twitter-RoBERTa (barbieri-etal-2020-tweeteval) 0.58 0.64 0.61
GPT-4o (openai2024systemcard) 0.67 0.82 0.74
@@
Table 1: Precision, recall, and F1 scores for the negative sentiment class across candidate models on our labelled reply set.
@@
Qwen 3 yang2025qwen3 achieved the strongest overall performance and is used as $s(\cdot)$ in all subsequent experiments. During error analysis, we found that most misclassifications occur in examples with low inter-annotator agreement. These instances often involved ambiguous language that led both assessors and models to varied interpretations.
6.4 Step 2 (Q2): Estimating PONOS from Post Text
With $s(\cdot)$ fixed, we compute PONOS from observed replies and then ask whether it can be accurately predicted at inference time (i.e. before replies are available).
Approach A: Direct PONOS Estimation
We train regression models to predict the measured PONOS value from the post text alone. This estimates reception without explicit reaction modelling and serves as a lightweight baseline suitable for low-latency settings.
Approach B: PONOS via Modelled Reactions
We also evaluate a reaction-modelling pipeline: a language model generates plausible replies to a post, and we apply $s(\cdot)$ to these generated replies to estimate PONOS. This directly instantiates CSF’s hypothesis that toxicity is reflected in likely community responses.
To separate the contributions of domain adaptation and reaction modelling, we evaluate two variants using LLaMA 3 70B llama2024herd: (i) reply prediction (RP) with PONOS inference via $s(\cdot)$, and (ii) domain-tuned reply prediction (DTRP).
6.4.1 Domain-Specific Pretraining Details
@@
For DTRP, we pretrained LLaMA 3 70B on r/BlackPeopleTwitter data to better capture community style and norms. We used the unsloth framework to load and fine-tune the base model in 4-bit precision. The model was adapted using LoRA with rank 64 for efficient parameter updates. Gradient checkpointing and rslora were employed to further reduce memory overhead. All training and inference were conducted using bfloat16 precision and FlashAttention 2 to maximize memory efficiency.
@@
Each training example consists of a Reddit comment and a prompt requesting five likely replies in the tone and style of r/BlackPeopleTwitter, separated by <|reply|>. The prompt template appears in Appendix E.
6.4.2 Step 3 (Q3): Comparing PONOS to Text-only Toxicity Scorers
We characterise how PONOS reception-based signal relates to intrinsic-toxicity instruments. To study divergence, we include intrinsic toxicity systems as reference instruments: the OpenAI Moderation API (omni-moderation-latest) (openai_moderation) and the Google Perspective API (jigsaw_perspective).
These systems estimate toxicity as a property of text and are therefore aligned with the perceived norm-violation component of our definition. In contrast, PONOS measures the reception or stress-disruption component via replies. We therefore focus on the relationship between these signals and on systematic regions of disagreement.
6.5 Results: Estimability and Construct Divergence
Estimating continuous PONOS.
Table 2 reports mean squared error (MSE) and mean absolute error (MAE) for continuous PONOS estimation. Two patterns emerge. First, strong text-only estimators (e.g. fine-tuned SentenceBERT) achieve competitive error, suggesting that some correlates of reception are recoverable from post text alone. Second, reaction modelling benefits substantially from domain adaptation: reply prediction without adaptation (RP) yields high error, while the domain-tuned reply prediction (DTRP) model attains the lowest error (MSE 0.042; MAE 0.152).
Together, these results support the claim that approximating community response can improve reception prediction when the generator is calibrated to community norms.
@@ 6.5 Results: Estimability and Construct Divergence
Table 2: Mean squared error (MSE) and mean absolute error (MAE) for PONOS prediction. FT = fine-tuned model.
Construct relationship analysis: PONOS vs. intrinsic instruments.
We quantify the association between reception-based measurement and intrinsic-toxicity scoring using large-scale comparisons (n=88,717n=88{,}717) between measured PONOS and two widely deployed toxicity detection instruments: the OpenAI Moderation API and the Google Perspective API. For OpenAI, we summarise each post by the maximum category score ("max moderation score") and additionally report correlations for individual categories. For the Perspective API, we use the continuous toxicity score. [^1] In all cases, we report Spearman’s $\rho$ (monotonic association) and Pearson’s $r$ (linear association), with 95% confidence intervals via Fisher’s $z$ approximation.
Both instruments show positive but modest association with reception. For OpenAI, PONOS correlates with the max moderation score at $\rho=0.203$ (95% CI [0.197, 0.210]) and $r=0.195$ (95% CI [0.189, 0.201]); among categories, harassment shows the strongest monotonic association ($\rho=0.207$). For the Perspective toxicity score, the association is $\rho=0.176$ (95% CI [0.170, 0.183]) and $r=0.173$ (95% CI [0.167, 0.179]). Together, these results support the construct claim: intrinsic-toxicity scoring and reception-based PONOS are related but non-equivalent signals.
Finally, to test whether these patterns depend on a particular instrument, we measure inter-tool agreement. OpenAI max moderation score and Perspective toxicity correlate strongly ($\rho=0.866$, 95% CI [0.864, 0.868]), indicating that intrinsic instruments form a coherent axis that is weakly aligned with the reception-based PONOS axis.
Figure 1: Two-axis view of intrinsic toxicity (OpenAI max moderation score; x-axis) versus reception-based toxicity (PONOS; y-axis). Each quadrant reports prevalence (count and percentage; n=88,717n=88{,}717) and includes illustrative masked snippets.
Quadrant breakdown.
We discretise both axes into high versus low using quantile thresholds and count posts in each quadrant. Using top-25% thresholds on PONOS and max moderation score (Figure 1), we find that off-diagonal disagreement is substantial. For the OpenAI max moderation score, 14,513 posts (16.4%) are high-PONOS/low-moderation and 14,158 (16.0%) are low-PONOS/high-moderation (32.3% off-diagonal total). The Perspective API shows the same pattern: 15,174 posts (17.1%) are high-PONOS/low-intrinsic and 14,887 (16.8%) are low-PONOS/high-intrinsic (33.9% off-diagonal).
This supports the view that reception and intrinsic toxicity are related but non-equivalent axes.
To interpret the disagreement regions, we analyse two extreme divergence sets constructed from posts where reception and intrinsic instruments disagree most strongly. We emphasise that these are not universal "model failures"; rather, they are diagnostic cases that make construct mismatch concrete.
Reception-high / instrument-low: missed norm violations.
These posts elicit consistently hostile, aggressive, or sarcastic replies (maximal negative reception) yet receive low intrinsic scores. On average, the OpenAI max moderation score is 0.090 and the Perspective API toxicity score is 0.155.
These cases often involve sarcasm, pragmatic antagonism, or contextually coded norm violations that are legible to the community but lack strong lexical markers.
Reception-low / instrument-high: overflagging benign dialect and in-group language.
Conversely, these posts receive no negative reception but are flagged as highly toxic by intrinsic scorers. The OpenAI max moderation score averages 0.757 and the Perspective API toxicity score averages 0.674.
The content is dominated by dialectal expressions, humour, and identity-referential language that is benign in-community but resembles out-group toxic markers in generic training data.
Takeaway.
Across both extremes, divergence is systematic: lexically "clean" antagonism can provoke strong negative reception while remaining low-scoring intrinsically, whereas culturally situated profanity and in-group language can be intrinsically flagged while eliciting no negative reception. This provides concrete validity evidence for CSF’s two-axis view: reception-based measurement is complementary to intrinsic toxicity, and helps avoid both missed harms and false alarms in dialect- and context-rich communities.
7 Conclusion
We reframed toxicity as a contextual, interactional phenomenon that arises when a communicative act is perceived to violate local norms and thereby induces stress or disruption in its audience.
We proposed the Contextual Stress Framework (CSF) and operationalised a reception-based signal, PONOS, as the proportion of negative reactions to an utterance. Empirically, we provided validity evidence for this reception axis: we selected an in-domain sentiment instrument for $s(\cdot)$, we showed that PONOS is estimable at inference time (with domain-tuned reaction modelling performing best), and demonstrated systematic construct divergence from intrinsic toxicity instruments. In particular, large and interpretable off-diagonal regions show that reception-based measurement is complementary to intrinsic toxicity, and can help avoid both missed harms and false alarms in dialect- and context-rich communities.
Taken together, these findings support CSF’s central claim: intrinsic-toxicity scores and reception-based measurement are related but non-equivalent. They form complementary axes for analysing harm in context.
Future work can extend this approach by developing richer reception signals beyond reply polarity, validating CSF across platforms and cultures, incorporating multimodal context, and studying interventions that use both axes to mitigate harm.